OpenStudy (anonymous):

Vectors A and B lie in an xy plane. A has magnitude 7.60 and angle 134° from the positive direction of the x axis. B has components Bx = 7.84 and By = -9.20. (Type 'undefined' for any answer if appropriate.) (a) What is 3A·B? (b) What is 4A 3B in unit-vector notation? k (c) What is 4A 3B in magnitude-angle notation with spherical coordinates (see Fig. 3-36)? Fig. 3-36 , θ = °, = ° (d) What is the angle between the directions of A and 4A 3B? (Hint: Think a bit before you resort to a calculation.) ° (e) What is A + 6.00 k in unit-vector notation? i + j + k (f) What is A

OpenStudy (anonymous):

I believe you may be missing some symbols for part b and c. (a) 3A = (3x7.60 cos(134 ), 3x7.60 sin(134)) = (-15.83, 16.40) B = (7.84, -9.20) 3A · B = -15.83 x 7.84 + 16.40 * -9.20 = -274.99